# Residual lifetime prediction for heterogeneous degradation data by Bayesian semi-parametric method

### v1

@@ 4.1 Conditional distributions for Gibbs sampling
    Conditional distribution of p given $(\bm{\beta},\mu_{\alpha},\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is represented in terms of distribution of $V_{h}^{*}$’s, $h=1,...,(N-1)$, where $p_{1}=V_{1}^{*}$ and $p_{h}=(1-V_{1}^{*})(1-V_{2}^{*})...(1-V_{h-1}^{*})V_{h}^{*}\hskip 2.84526pt,h=2,...,(N-1)$ and $p_{N}=1-\sum\limits_{h=1}^{N-1}p_{h}$ and the $V_{h}^{*}$ are independent and $V_{h}^{*}\sim Beta(1+r_{h},\gamma+\sum\limits_{l=k+1}^{N}r_{l})$, where $r_{h}=\sum\limits_{i=1}^{n}I_{K_{i}=h}$, $h=1,.......,N$.
@@
    Conditional distribution of ${\mu}_{h}$ given $(\bm{\beta},\alpha,\sigma_{\alpha}^{2},\bm{\mu}_{(-h)},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y}=\bm{y})$, for each $h\in[{K_{1}}^{*},.....,{K_{m}}^{*}]$ is $N(\mu_{h}^{*},{\sigma_{h}^{2}}^{*})$, where $\mu_{h}^{*}={\sigma_{h}^{2}}^{*}\times\sum\limits_{i:K_{i}=h}(X_{i}/{\sigma_{h}^{2}}+m_{\mu}/\sigma_{z})$, and ${\sigma_{h}^{2}}^{*}=\frac{{\sigma_{h}^{2}}{\sigma_{z}^{2}}}{r_{h}{\sigma_{z}^{2}}+{\sigma_{h}^{2}}}$. For $h\in\textbf{K}-[{K_{1}}^{*},.....,{K_{m}}^{*}]$ independently simulate from $N(m_{\mu},{\sigma_{z}^{2}})$.
@@
    Conditional distribution of $(\sigma_{z}^{2})^{-1}$ given $(\bm{\beta},\mu_{\alpha},\sigma_{\alpha}^{2},\bm{\mu},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is,
    $\mathcal{G}(\tau_{1}+\frac{N}{2},\tau_{2}+\sum\limits_{h=1}^{N}(\mu_{h}-m_{\mu})^{2})$.
  10. 10.
    Conditional distribution of $(\sigma_{h}^{2})^{-1}$ given $(\bm{\beta},\alpha,\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2}_{(-h)},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$, for each $h\in[{K_{1}}^{*},.....,{K_{m}}^{*}]$ is $\mathcal{G}(a+\frac{r_{h}}{2},b+\sum\limits_{i:K_{i}=h}(\beta_{i}-\mu_{h})^{2})$. For $h\in\textbf{K}-[{K_{1}}^{*},.....,{K_{m}}^{*}]$ independently simulate from $\mathcal{G}(a,b)$.
  11. 11.
    Conditional distribution of concentration parameter $\gamma$ given $(\bm{\beta},\mu_{\alpha},\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\bm{Y})$ is,
@@
We simulate total 200000 MCMC samples from the posterior distributions and discard initial 50000 samples as burn-in. We consider only the observations with lag sizes 50, to decrease the autocorrelation. The convergence of the generated Markov chains is validated by auto-correlation plot and trace plot. We compute the residual lifetime distribution of a unit based on the generated observations from posterior distribution of parameters. Next we discuss on generating observations from the derived residual lifetime distribution.
@@ 4.2 Transformation based Markov chain Monte Carlo (TMCMC)
Once we estimate the probability density function of residual lifetime distribution, our next goal is to predict residual lifetime for a unit functioning at some point $t_{k}$. One way to predict is simulating observation from the estimated distribution and a specific sample quantile of that generated samples can be used as predicted residual lifetime. MCMC techniques like Metropolis-Hastings (MH) helps to simulate observation from a given target distribution. However, several challenges arise for implementation of this algorithm. For example, a very large number of iterations (of the order of millions) are usually necessary for simulating observations from the target distribution. Another key issue for implementing MH algorithm is that it often leads to poor acceptance rates. To overcome these issues, we use a MCMC technique called Transformation based Markov chain Monte Carlo (TMCMC) algorithm introduced by Dutta and Bhattacharya [30] for simulating samples from the estimated residual lifetime distribution. This algorithm produces a rapidly mixing Markov chain and a better acceptance rate compared to the standard MH algorithms. For univariate case, TMCMC can be reduced to a MH algorithm with a specific proposal distribution, though in higher dimensions the proposal does not admit a mixture form and TMCMC cannot be a special case of the MH algorithm.
@@
Consider $S$ be the state space and suppose $T:S\times S^{\prime}\rightarrow S$ for some $S^{\prime}$ (may be a subset of $S$) is a differentiable transformation. TMCMC technique is based on constructing forward and backward transformations $T$ which are to be defined in such a way that the detailed balance and irreducibility hold for the Markov chain generated by this technique. If the current state is $x$, then the forward move is proposed by $x^{\prime}=x\epsilon$, where $\epsilon\in(0,1)$ is a simulated value from some arbitrary density of the form $g(\epsilon)I_{\epsilon}(0,1)$. The backward transformation $\frac{x^{\prime}}{\epsilon}$ is applied for moving back to $x$ from $x^{\prime}$. In general, for given $\epsilon$ and the current state $x$, forward transformation is denoted by $T(x,\epsilon)$, and the backward transformation by $T^{b}(x,\epsilon)$. One can observe that, the regions covered by the two transformations are disjoint. An important advantage associated with this algorithm is that whatever be the choice of the density $g(\epsilon)$, it cancels in the acceptance ratio of the proposed TMCMC algorithm. For more details see Dutta and Bhattacharya [30].
@@
The transformation $T$ is considered to be a differentiable transformation and the corresponding Jacobian is constructed as $J(x,\epsilon)=|\frac{\partial(T(x,\epsilon),\epsilon)}{\partial(x,\epsilon)}|$ which is non-zero almost everywhere. In this article we use the transformation, where $S=S^{\prime}=(0,\infty)$ and $T(x,\epsilon)=x\epsilon$. For all $x\in S$, $T^{b}(x,\epsilon)=\frac{x}{\epsilon}$; $A=(0,1)$. In this case Jacobian is $\epsilon$. Suppose $g$ is a density on $A$ where $A$ is a subset of $S^{\prime}$ such that $T(x,A)$ and $T^{b}(x,A)$ are disjoint and $10<p<$ 1, $\alpha(\cdot)$ be the acceptance ratio and $\pi(\cdot)$ be the target distribution. Consider an initial value $x_{0}$ and $x_{t}$ be the value that the chain takes at $t$th iteration, where $t\geq 1$. Then the MCMC algorithm based on transformation is constructed as follows.
@@
In this study the target probability density function is the density function as presented in equation (14). We have simulated 50000 MCMC iterates after discarding the initial 5000 iterations as burn-in. We take samples from lag size 10 from the rest of the samples and use the median of this as our predicted value of residual lifetime. Consider the predicted life as ${T_{ip}}$ and actual life as $T_{ia}$ for the $i$th individual unit. The prediction accuracy for $i$th individual unit is assessed by the relative absolute error $\frac{|T_{ip}-T_{ia}|}{T_{ia}}$. Figure 1 represents framework of our proposed Bayesian semi-parametric methodology which comprises of modeling degradation data to predicting residual lifetime of an individual unit.
Degradation data of $n$ units and a new unit with some degradation observations General Path model using $(n+1)$ units is built ( A mixed effect model) Dirichlet process mixture prior for the random effect distribution and parametric priors for fixed effect and variance of measurement errors is considered Simulate from the posterior distribution of parameters using Gibbs sampling as described above in Section 4.1 Use the posterior samples to compute the residual lifetime distribution of the new unit Use TMCMC algorithm to simulate observations from a derived distribution Compute median of the samples generated and use this as predicted value of residual lifetime and compute accuracy of prediction
Figure 1: Framework of proposed method
@@ 5 Simulation study
In this section we perform two simulation experiments to demonstrate the performance of the proposed Bayesian semi-parametric degradation modeling approach. We consider two and three component mixture distributions for random effects for two experiments. For the first case, we consider the true random effect distribution as $\beta\sim 0.6N(3,0.1)+0.4N(6,0.2)$. We consider true fixed effect $\alpha$ as 2 and the distribution of error represented as $\epsilon_{ij}\overset{\mathrm{i.i.d}}{\sim}N(0,0.5)$. The threshold value $D$ ts taken as 9. We generate degradation measures for 10 units and for each unit the degradation is measured in the time interval $(0,1)$ at equally spaced times $\frac{1}{10}$, until the degradation measure crosses the threshold value $D=9$, that is if for some unit degradation measures crosses threshold before the maximum time which is 1 in this case, we stop observing data and which implies that $n_{i}\leq 10$ for $i=1,...,10$. We assume the true linear model $Y_{ij}=\alpha+\beta_{i}t_{ij}+\epsilon_{ij},\hskip 5.69054pti=1,....,10,j=1,....,n_{i}$. For the second case we consider the true random effect distribution with mixture of 3 normal distribution represented as $\beta\sim 0.4N(2,0.1)+0.3N(3,0.15)+0.3N(6,0.12)$. Here also we take fixed effect $\alpha$ as 2 and $\epsilon_{ij}\overset{\mathrm{i.i.d}}{\sim}N(0,0.6)$ and $D$ as 9. We generate degradation measures for 30 units and for each unit the degradation is measured according to the plan followed for case 1.
@@
We run the simulation procedure to generate degradation path from the above-mentioned models in both the cases. We generate observations using the sample interval of 0.001 and obtain the time when the path crosses the predefined threshold for the first time. In this work we predict residual lifetime of each unit at the last time point when the observed degradation measurement has not crossed threshold value. We evaluate the prediction accuracy of the residual life using the error criteria introduced in the Section 4.2. The proposed Bayesian semi-parametric method is compared with a Bayesian parametric method. For Bayesian parametric method, we consider $\beta_{i}\sim N(\mu_{\beta},\sigma_{\beta}^{2})$, and $\alpha\sim N(\mu_{\alpha},\sigma_{\alpha}^{2})$ and $(\sigma_{\epsilon}^{2})^{-1}\sim G(a_{\epsilon},b_{\epsilon})$. The prior specification is: $\mu_{\beta}\sim N(m_{\mu},\sigma_{z}^{2})$, $({\sigma_{z}^{2}})^{-1}\sim\mathcal{G}(\tau_{1},\tau_{2})$ and $(\sigma_{\beta}^{2})^{-1}\sim G(a,b)$ where $a=\sqrt{v_{\beta}},b=\sqrt[3]{v_{\beta}}$. We choose $\mu_{\alpha}\sim N(0,10^{3})$ and $(\sigma_{\alpha}^{2})^{-1}\sim G(0.001,0.001)$. We implement a similar procedure discussed in section 4 to choose values for $m_{\mu}$ and $v_{\beta}$. $(\tau_{1},\tau_{2})=(0.001,0.001)$ and $(a_{\epsilon},b_{\epsilon})=(0.001,0.001)$ is considered for this case. The degradation path simulated from the two cases are represented in Figure 2.
Figure 2: Degradation paths generated from two different cases: (a) Case 1 (b) Case 2.

### v2

@@ 4.1 Conditional distributions for Gibbs sampling
    a) Conditional distribution of $(\sigma_{z}^{2})^{-1}$ given $(\bm{\beta},\bm{\mu},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is, $\mathcal{G}(\tau_{1}+\frac{N}{2},\tau_{2}+\sum\limits_{h=1}^{N}(\mu_{h}-m_{\mu})^{2})$, for gamma prior on $(\sigma_{z}^{2})^{-1}$.
    b) Conditional distribution of $\sigma_{z}$ given $(\bm{\beta},\bm{\mu},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is denoted as $f(\sigma_{z}|.)$, where $f(\sigma_{z}|.)\propto(\frac{1}{\sigma_{z}})^{N}\exp\Big[-\frac{1}{2\sigma_{z}^{2}}\sum\limits_{h=1}^{N}(\mu_{h}-m_{\mu})^{2}\Big]\times\Big(1+\frac{\sigma_{z}^{2}}{A_{2}}\Big)^{-1}$ for half-cauchy prior for $\sigma_{z}$.
@@
    a) Conditional distribution of $(\sigma_{h}^{2})^{-1}$ given $(\bm{\beta},\alpha,\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2}_{(-h)},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$, for each $h\in[{K_{1}}^{*},\dots,{K_{m}}^{*}]$ is, is $\mathcal{G}(a+\frac{r_{h}}{2},b+\sum\limits_{i:K_{i}=h}(\beta_{i}-\mu_{h})^{2})$ and for $h\in\textbf{K}-[{K_{1}}^{*},\dots,{K_{m}}^{*}]$ independently simulate from $\mathcal{G}(a,b)$ for Gamma prior of $(\sigma_{h}^{2})^{-1}$.
    b) Conditional distribution of $\sigma_{h}$ given $(\bm{\beta},\alpha,\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2}_{(-h)},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$, for each $h\in[{K_{1}}^{*},\dots,{K_{m}}^{*}]$ is denoted as $f(\sigma_{h}|.)$, where $f(\sigma_{h}|.)\propto(\frac{1}{\sigma_{h}})^{r_{h}}\exp\Big[-\frac{1}{2\sigma_{h}^{2}}\sum\limits_{i:K_{i}=h}(\beta_{i}-\mu_{h})^{2}\Big]\times\Big(1+\frac{\sigma_{h}^{2}}{A_{1}}\Big)^{-1}$ and for $h\in\textbf{K}-[{K_{1}}^{*},\dots,{K_{m}}^{*}]$ independently simulate from $\text{Half-Cauchy}(A_{1})$ for half-cauchy prior of $\sigma_{h}$.
@@
    Conditional distribution of concentration parameter $\gamma$ given $(\bm{\beta},\alpha,\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\bm{Y})$ is,
@@
For simulating observations from the conditional distribution given by 7-b) and 8-b) we use Metropolis-Hastings algorithm, where the proposal distribution is taken as gamma distribution with shape parameter 1 and rate parameter as the last iterated value. So, suppose that we start with initial value $\sigma_{z}^{0}$, then for the $(i+1)$th step in Metropolis-Hastings, we draw values from $\mathcal{G}(1,\sigma_{z}^{i})$, where $\sigma_{z}^{i}$ is the value obtained in the $i$th step. For semi-parametric method we take truncation number $N$ as the number of units in the study for each case. We simulate total 50000 MCMC samples from the posterior distributions for each scenario and discard initial 5000 samples as burn-in. We consider only the observations with lag sizes 50, to decrease the autocorrelation. The convergence of the generated Markov chains is validated by auto-correlation plot and trace plot. We compute the residual lifetime distribution of a unit based on the generated observations from posterior distribution of parameters. To make the residual lifetime distribution valid as discussed in section 3, if there is any negative value generated for the conditional distribution of the random effects by the Gibbs sampling method, we discard them and take only the samples in $(0,\infty)$. Next we discuss on generating observations from the derived residual lifetime distribution given by equation (12) and (16).
@@ 4.2 Transformation based Markov chain Monte Carlo (TMCMC)
Once we estimate the probability density function of residual lifetime distribution, our next goal is to predict residual lifetime for a unit functioning at some point $t_{k}$. One way to predict is simulating observation from the estimated distribution and a specific sample quantile of that generated samples can be used as predicted residual lifetime. MCMC techniques like Metropolis-Hastings (MH) helps to simulate observation from a given target distribution. However, several challenges arise for implementation of this algorithm. For example, a very large number of iterations (of the order of millions) are usually necessary for simulating observations from the target distribution. Another key issue for implementing MH algorithm is that it often leads to poor acceptance rates. To overcome these issues, we use a MCMC technique called Transformation based Markov chain Monte Carlo (TMCMC) algorithm introduced by Dutta and Bhattacharya [46] for simulating samples from the estimated residual lifetime distribution. This algorithm use simple deterministic bijective transformations of a low-dimensional random variable from some arbitrary distribution to update parameters simultaneously within a block. This algorithm produces a rapidly mixing Markov chain and a better acceptance rate compared to the standard MH algorithms ( see [47], [48]). For univariate case, TMCMC can be reduced to a MH algorithm with a specific proposal distribution, though in higher dimensions the proposal does not admit a mixture form and TMCMC cannot be a special case of the MH algorithm.
@@
Consider $S$ be the state space and suppose $T:S\times S^{\prime}\rightarrow S$ for some $S^{\prime}$ (may be a subset of $S$) is a differentiable transformation. TMCMC technique is based on constructing forward and backward transformations $T$ which are to be defined in such a way that the detailed balance and irreducibility hold for the Markov chain generated by this technique. Consider a transformation, where the current state is $x$, then the forward move is proposed by $x^{\prime}=x\epsilon$, where $\epsilon\in(0,1)$ is a simulated value from some arbitrary density of the form $g(\epsilon)I_{\epsilon}(0,1)$. The backward transformation $\frac{x^{\prime}}{\epsilon}$ is applied for moving back to $x$ from $x^{\prime}$. In general, for given $\epsilon$ and the current state $x$, forward transformation is denoted by $T(x,\epsilon)$, and the backward transformation by $T^{b}(x,\epsilon)$. One can observe that, the regions covered by the two transformations are disjoint. An important advantage associated with this algorithm is that whatever be the choice of the density $g(\epsilon)$, it cancels in the acceptance ratio of the proposed TMCMC algorithm. For more details see Dutta and Bhattacharya [46].
The transformation $T$ is considered to be a differentiable transformation and the corresponding Jacobian is constructed as $J(x,\epsilon)=|\frac{\partial(T(x,\epsilon),\epsilon)}{\partial(x,\epsilon)}|$ which is non-zero almost everywhere. In this article, we simulate from the residual lifetime having density function given in equation (13) and (17). Under $f_{T_{new}|\bm{Y}}^{approx}(t)$, residual lifetime can take negative values whereas, residual lifetime has support on $(0,\infty)$ with respect to $f_{T_{new}|\bm{Y},T_{new}>0}^{approx}(t)$. We use the following transformations,
  1. a)
    For simulating observations with respect to pdf given in equation (13), we take $S=S^{\prime}=(-\infty,\infty)$ and $T(x,\epsilon)=x+\epsilon$. For all $x\in S$, $T^{b}(x,\epsilon)=x-\epsilon$; $A=(0,\infty)$, with Jacobian is 1.
  2. b)
    For simulating observations with respect to pdf given in equation (17), we consider $S=S^{\prime}=(0,\infty)$ and $T(x,\epsilon)=x\epsilon$. For all $x\in S$, $T^{b}(x,\epsilon)=\frac{x}{\epsilon}$; $A=(0,1)$, with Jacobian is $\epsilon$.
@@
Suppose $g$ is a density on $A$ where $A$ is a subset of $S^{\prime}$ such that $T(x,A)$ and $T^{b}(x,A)$ are disjoint and $0<p<1$, $\alpha(\cdot)$ be the acceptance ratio and $\pi(\cdot)$ be the target distribution. Consider an initial value $x_{0}$ and $x_{t}$ be the value that the chain takes at $t$th iteration, where $t\geq 1$. Then the MCMC algorithm based on transformation is constructed as follows.
@@
We have simulated 10000 MCMC iterates after discarding the initial 1000 iterations as burn-in. We take samples from lag size 10 from the rest of the samples and use the median of this as our predicted value of residual lifetime. For this procedure, we take initial value as $x_{0}=1$ and $p=\frac{1}{2}$ for each case.
5 Prediction evaluation criteria
Suppose there are $n_{1}$ many units for which residual lifetime is to be predicted. Consider the predicted life as $T_{ip}$} and actual life as $T_{ia}$ for the $i$th individual unit. The prediction accuracy for $i$th individual unit is assessed by the root mean squared error (RMSE) and mean absolute error (MAE) which are given as,
  1. 1.
    RMSE = $\sqrt{\frac{1}{n_{1}}\sum\limits_{i=1}^{n_{1}}(T_{ip}-T_{ia})^{2}}$
  2. 2.
    MAE = $\frac{1}{n_{1}}\sum\limits_{i=1}^{n_{1}}|T_{ip}-T_{ia}|$
We evaluate both the proposed Bayesian semi-parametric method and parametric method with respect to these error criteria. Lower values of RMSE and MAE indicates better prediction.
@@ 6 Simulation study
In this section we perform different simulation experiments to demonstrate the performance of the proposed Bayesian semi-parametric degradation modeling approach. We consider three component mixture of gamma and Weibull distributions for random effects for first two experiments. For the last three experiments, we consider two, three and five component mixture of normal distributions. For each case, we consider true fixed effect $\alpha$ as 2. The threshold value $D$ and the distribution of error measurement and random effect used for the study are given in Table 3. For each case, we generate degradation measures for $n=10,30,50$ units. Degradation observations are generated in the time interval $[0,1]$ at equally spaced times of length $\frac{1}{10}$ and $\frac{1}{30}$ respectively. We stop generating observations if degradation measure crosses the respective threshold value $D$ before the maximum time or till the last time point. So, maximum number of observations, denoted as $m=\max\{n_{i}\}$ per unit is 11 and 31 respectively for $i=1,\dots,n$. Suppose, for some unit degradation measures crosses threshold before the maximum time which is 1 in this case, we stop observing data. We generate observations assuming the true degradation model as, $Y_{ij}=\alpha+\beta_{i}t_{ij}+\epsilon_{ij},\hskip 5.69054ptj=1,\dots,n_{i},i=1,\dots,n$.
Table 3: Scenarios for Simulation Study
Case         Distributions and Threshold value
1 $\beta_{i}\overset{\mathrm{i.i.d}}{\sim}0.4\mathcal{G}(40,20)+0.25\mathcal{G}(70,20)+0.35\mathcal{G}(100,20)$, $\epsilon_{ij}\overset{\mathrm{i.i.d}}{\sim}\mathcal{N}(0,0.7)$, $D=11$.
2 $\beta_{i}\overset{\mathrm{i.i.d}}{\sim}0.4\text{Weibull}(20,2)+0.25\text{Weibull}(40,4)+0.35\text{Weibull}(80,8)$, $\epsilon_{ij}\overset{\mathrm{i.i.d}}{\sim}\mathcal{N}(0,0.7)$ , $D=12$.
3 $\beta_{i}\overset{\mathrm{i.i.d}}{\sim}0.6\mathcal{N}(3,0.1)+0.4\mathcal{N}(6,0.2)$, $\epsilon_{ij}\overset{\mathrm{i.i.d}}{\sim}\mathcal{N}(0,0.5)$, $D=9$.
4 $\beta_{i}\overset{\mathrm{i.i.d}}{\sim}0.4\mathcal{N}(2,0.1)+0.3\mathcal{N}(4,0.15)+0.3\mathcal{N}(6,0.12)$, $\epsilon_{ij}\overset{\mathrm{i.i.d}}{\sim}\mathcal{N}(0,0.6)$, $D=9$.
5 $\beta_{i}\overset{\mathrm{i.i.d}}{\sim}0.3\mathcal{N}(2,0.1)+0.2\mathcal{N}(4,0.15)+0.2\mathcal{N}(6,0.12)+0.15\mathcal{N}(8,0.15)+0.15\mathcal{N}(10,0.1)$, $\epsilon_{ij}\overset{\mathrm{i.i.d}}{\sim}\mathcal{N}(0,0.7)$, $D=13$.
@@
For gamma mixture, we consider parametrization of the form $\mathcal{G}(a_{1},b_{1})$ where shape parameter is $a_{1}$ and rate parameter is $b_{1}$ and for Weibull mixture, where each component is denoted as $\text{Weibull}(a_{2},b_{2})$ has shape parameter $a_{2}$ and scale parameter $b_{2}$. For each combination of $n$ and $m$, we conduct the analysis with respect to each priors described in Tables 1 and 2. The goal of this article is to predict residual lifetime of each unit, so it is required to compare the predicted value with the true value. To obtain the true value of residual lifetime, we run the simulation procedure to generate degradation path from the above-mentioned models. We generate observations using the sample interval of length 0.001 and obtain the time when the degradation path crosses the predefined respective threshold for the first time. We consider this as true value to check the prediction accuracy of the proposed methods. We predict residual lifetime of each unit at the last time point when the degradation measurement has not crossed threshold value $D$. We evaluate the prediction accuracy of the residual life using the error criteria introduced in Section 4.2. We compare both distribution functions introduced in equations (12) and (16) for residual lifetime prediction. Suppose $M_{1}$ denotes the method for residual lifetime prediction with constraint $T_{new}>0$ and $M_{2}$ without the constraint. The proposed Bayesian semi-parametric method is compared with a Bayesian parametric method. We present analysis for residual lifetime distribution based on prior 2 for each case. For each case, Kolmogorov-Smirnov (KS) statistic is calculated using the generated samples with respect to the residual lifetime distribution given in equation (15) to measure the distance between true and approximated distribution. The estimate for parameters and residual lifetime obtained by both the parametric and semi-parametric methods are rounded off to three decimal places for each cases. The degradation path simulated from the different cases are represented in Figure 1.
(a) Case 1
(b) Case 2
(c) Case 3
(d) Case 4
(e) Case 5
Figure 1: Degradation paths generated with $n=10$ and $m=31$.